# CS-Bench: A Comprehensive Benchmark for Large Language Models towards Computer Science Mastery

### v1

@@ Abstract
Computer Science (CS) stands as a testament to the intricacies of human intelligence, profoundly advancing the development of artificial intelligence and modern society. However, the current community of large language models (LLMs) overly focuses on benchmarks for analyzing specific foundational skills (e.g. mathematics and code generation), neglecting an all-round evaluation of the computer science field. To bridge this gap, we introduce CS-Bench, the first bilingual (Chinese-English) benchmark dedicated to evaluating the performance of LLMs in computer science. CS-Bench comprises approximately 5K meticulously curated test samples, covering 26 subfields across 4 key areas of computer science, encompassing various task forms and divisions of knowledge and reasoning. Utilizing CS-Bench, we conduct a comprehensive evaluation of over 30 mainstream LLMs, revealing the relationship between CS performance and model scales. We also quantitatively analyze the reasons for failures in existing LLMs and highlight directions for improvements, including knowledge supplementation and CS-specific reasoning. Further cross-capability experiments show a high correlation between LLMs’ capabilities in computer science and their abilities in mathematics and coding. Moreover, expert LLMs specialized in mathematics and coding also demonstrate strong performances in several CS subfields. Looking ahead, we envision CS-Bench serving as a cornerstone for LLM applications in the CS field and paving new avenues in assessing LLMs’ diverse reasoning capabilities. The CS-Bench data and evaluation code are available at https://github.com/csbench/csbench.
@@ 1 Introduction
Serving as the cornerstone of the modern information revolution, the significance of computer science (CS) extends from the early days of electronic computers to today’s advancements in artificial intelligence (AI) [1, 2]. As a new milestone in AI, large language models (LLMs) [3, 4] represented by ChatGPT [5] and GPT-4 [6] are not limited to the natural language processing (NLP) community, showing vast potential in fields including education, industry, and science [7, 8, 9, 10, 11, 12, 13]. However, enabling LLMs to effectively utilize computer science knowledge and serve humanity more efficiently is one of the key challenges on the path to the future intelligent era [14, 15, 16].
Understanding the performance of LLMs in computer science is fundamental to the research and application of LLMs within the field. Despite studies like MMLU and C-Eval [17, 18, 19, 20, 21] covering a wide range of fields including CS, their broad scope implies that CS is merely a component within the multiple categories of science and engineering, overlooking the importance of thoroughly evaluating the CS field. Moreover, such evaluation result can further guide the development of LLMs, offering practical insights to advance the corresponding capabilities. Recently, a series of studies have devoted on actively assessing and analyzing the capabilities of LLMs in mathematics, coding, and logical reasoning [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. Unfortunately, efforts on LLMs in cross-capability evaluation is quite scarce. Considering the intersection of computer science with coding, mathematics, and reasoning abilities, we have grounds to believe that cross-capability research and analysis in CS can effectively propel the comprehensive development of the LLM community. Here, we are particularly interested in two research questions for evaluating LLMs’ proficiency in computer science field:
@@
As the bedrock for exploration, we first propose CS-Bench, the first benchmark dedicated to evaluating the performance of LLMs in the field of computer science. CS-Bench features high-quality, diverse task forms, varying capacities, and bilingual evaluation. Firstly, CS-Bench comprises approximately 5,000 carefully curated test items spanning 26 sections across 4 key CS domains. Diverging from conventional benchmarks consisting solely of multiple-choice (MC) questions [17, 18, 20], CS-Bench includes 4 tasks: multiple-choice, assertion, fill-in-the-blank (FITB), and open-ended, to better simulate real-world scenarios and assess the robustness of LLMs to different task formats. In addition to knowledge-type questions assessing LLMs’ mastery of CS knowledge, reasoning-type questions further evaluate LLMs’ ability to apply CS knowledge for reasoning. Lastly, by supporting bilingual evaluation in Chinese and English, CS-Bench enables the appraisal of LLMs’ adeptness in addressing CS challenges across different language contexts.
@@
In response to RQ1, we evaluate over 30 mainstream LLMs on CS-Bench. Our main findings are: (1) CS-Bench effectively differentiates the capabilities of LLMs in the CS field while also posing significant challenges to the best-performing GPT-4/ GPT-4o. (2) LLMs exhibit a consistent logarithmic growth pattern in scale and a linear growth pattern in scores on the CS-Bench. By establishing the scale-score fitting function, smaller models can be used to predict and guide the development of larger-scale models. (3) Further error type analysis indicates that the primary reason for the limited performance of LLMs is the lack of domain knowledge, and the CS-specific reasoning is difficult to achieve merely by enhancing general reasoning abilities, necessitating targeted reinforcement.
@@
  • We introduce CS-Bench, the first benchmark dedicated to evaluate the performance of LLMs in the field of computer science. CS-Bench supports both Chinese and English, covers four key areas with 26 subfields, and includes a diverse range of task formats.
@@ 2.1 Design Principle
The objective of CS-Bench is to robustly assess the knowledge and reasoning capabilities of LLMs in different linguistic contexts within the field of computer science. To this end, our benchmark adheres to the following guidelines: (1) Coverage of key domains: it covers key areas of CS with finer subfields for specificity. (2) Diverse task forms: questions vary in format to simulate diverse real-world user queries. (3) CS-specific reasoning: it evaluates CS logical and arithmetic reasoning in addition to CS knowledge. (4) Multilinguality support: it supports assesses LLMs’ performance in different language environments. Based on these criteria, CS-Bench focuses on bilingual evaluation in Chinese and English, covering four domains: Data Structure and Algorithm (DSA), Computer Organization (CO), Computer Network (CN), and Operating System (OS). Twenty-six fine-grained subfields, diverse task forms, and divisions of knowledge and reasoning are further developed to enrich the dimensions of assessment and simulate real-world scenarios.
@@ 2.2 Data Collection
Table 1: Comparison of perplexity (PPL) across evaluation datasets. The PPL of English and Chinese datasets is calculated on Llama2-7B-base and Qwen1.5-7B-base, respectively. “MC” denotes multiple-choice, and “ALL” denotes all tasks.
@@
TruthfulQA (MC) [34] 7.73 C-Eval [18] 11.47
MMLU [35] 9.54 CMMLU [20] 13.62
@@
Diverse data sources are key to achieving the sample diversity of CS-Bench. Our raw data originates from three sources: (1) Computer science-related questions obtained from publicly available online channels, such as professional exams and practice tests111e.g., https://github.com/CodePanda66/CSPostgraduate-408. (2)Knowledge-type questions obtained through the initial manual extraction and subsequent adaptation of blog articles from various computer-related websites222e.g., https://www.wikipedia.org/,https://www.cnblogs.com/, https://www.csdn.net/. (3) Construction of teaching materials and examination papers authorized by the authors’ institutions. The latter two categories constitute the vast majority (over 70%) of the data, and these data are not directly exposed on the internet, effectively reducing the likelihood of LLMs encountering these questions during pre-training. We compare the perplexity [36] of models on CS-Bench and several prominent evaluation datasets in Table 1. In both English and Chinese, the perplexity of CS-Bench is comparable to or even higher than that of other datasets, further indicating the high quality of CS-Bench samples and the rarity of data leakage instances.
@@
The data processing relies on a team composed of five members, each holding a bachelor’s degree in computer science and receiving appropriate compensation. Initially, we parse questions and answers for each sample from the data sources either automatically or manually. Subsequently, we manually label questions with knowledge-type or reasoning-type tags depending on whether in-depth reasoning and calculation are required. For reasoning-type questions, we attempt to collect explanations from the data sources whenever possible; otherwise, we handle them through cross-annotation and verification among team members. We first construct Chinese data, then translate it into English using GPT-4, supplemented by manual checks, to create English data. Finally, we conduct thorough manual checks on the entire dataset to ensure quality. As this benchmark pertains to objective knowledge and reasoning in the field of computer science, the annotation content is not influenced by regional or cultural differences among annotators.
@@
CS-Bench is an evaluation benchmark supporting bilingual assessment, encompassing a total of 26 subfields across 4 domains, with a cumulative total of 4838 samples. These samples encompass various task formats including multiple-choice, assertion, fill-in-the-blank, and open-ended questions. Besides, CS-Bench assesses both knowledge-type and higher-order reasoning-type questions, with each reasoning question accompanied by an explanation. To validate the effectiveness of models, we randomly sample 10% of the data for validation, using the remaining 90% for testing. The statistics of CS-Bench are shown in Figure 1, with detailed exposition provided in Appendix C.
@@ 3.1 Experimental Setup
Due to the diverse task formats in CS-Bench, we first design question templates for each task type. For comprehension tasks (MC and Assertion), we use regex to match LLM’s predictions and then calculate their accuracy against the ground-truth answers. For generation tasks (FITB and Open-ended), due to the diversity of ground-truth answers, we score LLM’s predictions by GPT-4 using standard answers in CS-Bench as references. In detail, FITB questions are scored as either 0 or 1, while the score range for Open-ended questions is 1-10, which is then linearly mapped to a range of 0.1 to 1. Finally, scores are weighted based on the quantity of each type to derive the ultimate overall score. It is worth emphasizing that while employing GPT-4 for scoring generation tasks may introduce a certain threshold for evaluation, its primary purpose is to simulate diverse task formats in real-world scenarios. Therefore, we encourage isolating comprehension tasks from CS-Bench to facilitate automatic evaluation with no need for GPT-4. We provide the details of the evaluation setup in Appendix D, where we also verify the validity of GPT-4 scoring through its consistency with manually scored results.
@@
We evaluate nearly 30 models in different sizes from 12 model families. For open-source models, we selected Gemma-2B/7B [37], Llama2-7B/13B/70B [38], Llama3-8B/70B [39], ChatGLM3-6B [40], Baichuan2 (v2.0)-7B/13B [41], InternLM2-7B/20B [42] , Qwen1.5-4B/7B/14B/72B/110B [43], Mistral-7B (v0.2) [44], Mixtral-8$\times$7B (v0.1) [45], and DeepSeekLLM-7B/67B [46]. For closed-source commercial models, we utilized PaLM-2 (palm-2-chat-bison) [47], Claude-2.1 [48], Claude-3 (opus) [49], as well as GPT-3.5, GPT-4 (0125 version) [50] and GPT-4o [6]. To ensure the instruction-following abilities, we employ the official chat or instruction-tuned versions for all models. Details on these models are provided in Appendix D.4.
@@ 3.2 Main Results
Table 2 presents the overall results of all foundation models directly answering questions under the zero-shot setting  [^3] Due to space constraints, the results and analysis on CS-Bench (CN) are provided in Appendix E.3.. In summary, the overall scores of models range from 39.86% to 72.29%, demonstrating CS-Bench’s effectiveness in distinguishing between the abilities of various models in the field of CS while also posing significant challenges to the best-performing existing models. Subsequently, we conduct a comprehensive analysis of the experimental results from various aspects including Foundation Models, Knowledge & Reasoning, Domains, and Task Formats.
@@
Firstly, the closed-source models GPT-4/GPT-4o represent the highest standard on CS-Bench, being the only two models exceeding 70% proficiency. Secondly, the disparity between the leading open-source and closed-source models is not significant. Notably, premier open-source models such as Qwen1.5-110B and Llama3-70B have surpassed previously strong closed-source models like GPT-3.5 and Claude-2.1, drawing close to Claude-3 in performance. Thirdly, newer models demonstrate significant improvements compared to earlier versions. For example, among models with scales below 10B, Llama3-8B performs the best, rivaling previous much larger-scale models and even surpassing Llama2-70B, indicating significant potential for compression in model parameters [51]. Lastly, while performance variations exist among models of different families at the same scale, models within the same family continue to improve with increasing scale on CS-Bench (see detailed scale analysis in Section 3.3).
@@
Overall, all models perform worse on reasoning (average 44.63%) compared to knowledge scores (average 60.52%), indicating that reasoning poses a greater challenge to LLMs compared to knowledge. As shown in Figure 3, there is a strong positive correlation between reasoning scores and knowledge scores. However, this correlation is not absolute. For instance, PaLM-2 has a higher knowledge score but a lower reasoning score compared to Claude-2.1, showing PaLM-2’s weakness in applying knowledge. Furthermore, more powerful LLMs demonstrate a stronger ability to use knowledge for reasoning compared to weaker LLMs. This is reflected in the much lower reasoning scores of weaker models relative to their knowledge scores. However, as the model’s capability increases, the growth in reasoning scores is more pronounced than that of knowledge scores, gradually bridging the gap between knowledge and reasoning abilities.
@@
As shown in Figure 3 (b) and Table 13, given the varying initial random scores, LLMs generally performs best on Assertion questions (average 63.11% across all models), followed by MC questions (average 54.92%), Open-ended questions (average 49.1%), and performs worst on FITB questions (average 41%). However, the variation in task format sensitivity is highly pronounced in weaker models, while stronger models can mitigate the disparities caused by different task formats, exhibiting robustness. For instance, Llama2-7B scores only 26.19% on Open-ended reasoning but 60.61% on Assertion reasoning, whereas GPT-4 scores comparably on both Open-ended reasoning (68.94%) and Assertion reasoning (67.68%).

### v2

@@ Abstract
Large language models (LLMs) have demonstrated significant potential in advancing various fields of research and society. However, the current community of LLMs overly focuses on benchmarks for analyzing specific foundational skills (e.g. mathematics and code generation), neglecting an all-round evaluation of the computer science field. To bridge this gap, we introduce CS-Bench, the first multilingual (English, Chinese, French, German) benchmark dedicated to evaluating the performance of LLMs in computer science. CS-Bench comprises approximately 10K meticulously curated test samples, covering 26 subfields across 4 key areas of computer science, encompassing various task forms and divisions of knowledge and reasoning. Utilizing CS-Bench, we conduct a comprehensive evaluation of over 30 mainstream LLMs, revealing the relationship between CS performance and model scales. We also quantitatively analyze the reasons for failures in existing LLMs and highlight directions for improvements, including knowledge supplementation and CS-specific reasoning. Further cross-capability experiments show a high correlation between LLMs’ capabilities in computer science and their abilities in mathematics and coding. Moreover, expert LLMs specialized in mathematics and coding also demonstrate strong performances in several CS subfields. Looking ahead, we envision CS-Bench serving as a cornerstone for LLM applications in the CS field and paving new avenues in assessing LLMs’ diverse reasoning capabilities. Our project homepage is available at https://csbench.github.io/.
@@ 1 Introduction
Large language models (LLMs), exemplified by ChatGPT (OpenAI, 2022), have emerged as a significant milestone in artificial intelligence (AI), demonstrating potential far beyond natural language processing (NLP) (Zhao et al., 2023; Chang et al., 2024). These models are increasingly impacting diverse fields including education, industry, and scientific research (Guha et al., 2023; Guo et al., 2023; Xiao et al., 2023; Huang et al., 2023a; Zhou et al., 2023a; Zhao et al., 2024; Zhang et al., 2024a). As computer science (CS) continues to be the cornerstone of modern information technology, from the advent of electronic computers to today’s AI advancements (Denning, 2000; Campbell-Kelly et al., 2023), a key challenge lies in enabling LLMs to better utilize CS knowledge to serve humanity in the evolving intelligent era (Donadel et al., 2024; Murtuza, 2024; Marques et al., 2024).
Understanding the performance of LLMs in computer science is fundamental to the research and application of LLMs within the field. Despite studies like MMLU and C-Eval (Hendrycks et al., 2021a; Huang et al., 2023b; Liu et al., 2023a; Li et al., 2024; Gu et al., 2024) covering a wide range of fields including CS, their broad scope implies that CS is merely a component within the multiple categories of science and engineering, overlooking the importance of thoroughly evaluating the CS field. Moreover, such evaluation result can further guide the development of LLMs, offering practical insights to advance the corresponding capabilities. Recently, a series of studies have devoted on actively assessing and analyzing the capabilities of LLMs in mathematics, coding, and logical reasoning (Frieder et al., 2023; Collins et al., 2023; Wu et al., 2023; Yuan et al., 2023b; Dong et al., 2024; Liu et al., 2024; Zhang et al., 2024b; Lin et al., 2024; Liu et al., 2023b; Saparov et al., 2023; Xu et al., 2023; Wu et al., 2024). Unfortunately, efforts on LLMs in cross-capability evaluation is quite scarce. Considering the intersection of computer science with coding, mathematics, and reasoning abilities, we have grounds to believe that cross-capability research and analysis in CS can effectively propel the comprehensive development of the LLM community. Here, we are particularly interested in two research questions for evaluating LLMs’ proficiency in computer science field:
@@
As the bedrock for exploration, we introduce CS-Bench, the first benchmark dedicated to evaluating the performance of LLMs in computer science. CS-Bench features high-quality, diverse task forms, varying capacities, and quadrilingual evaluation. Firstly, CS-Bench comprises approximately 10K carefully curated test items spanning 26 sections across 4 key CS domains. Unlike conventional benchmarks consisting solely of multiple-choice (MC) questions (Hendrycks et al., 2021a; Huang et al., 2023b; Li et al., 2024), CS-Bench includes 4 tasks: multiple-choice, assertion, fill-in-the-blank (FITB), and open-ended, to better simulate real-world scenarios and assess the robustness of LLMs to different task formats. In addition to knowledge-type questions assessing LLMs’ mastery of CS knowledge, reasoning-type questions further evaluate LLMs’ ability to apply CS knowledge for reasoning. Lastly, by supporting evaluation in English, Chinese, French, and German, CS-Bench allows assessment of LLMs’ ability to tackle CS challenges in various language contexts.
@@
In response to RQ1, we evaluate over 30 mainstream LLMs on CS-Bench. Our main findings are: (1) CS-Bench effectively differentiates the capabilities of LLMs in the CS field while also posing challenges to the best-performing GPT-4o/OpenAI-o1. (2) LLMs exhibit a consistent logarithmic growth pattern in scale and a linear growth pattern in scores on the CS-Bench. By establishing the scale-score fitting function, smaller models can be used to predict and guide the development of larger-scale models. (3) Further error type analysis indicates that the primary reason for the limited performance of LLMs is the lack of domain knowledge, and the CS-specific reasoning is difficult to achieve merely by enhancing general reasoning abilities, necessitating targeted reinforcement.
@@
  • We introduce CS-Bench, the first benchmark dedicated to evaluate the performance of LLMs in the field of computer science. CS-Bench supports four languages, covers four key areas with 26 subfields, and includes a diverse range of task formats.
@@ 2.1 Design Principle
The objective of CS-Bench is to robustly assess the knowledge and reasoning capabilities of LLMs in different linguistic contexts within the field of computer science. To this end, our benchmark adheres to the following guidelines: (1) Coverage of key domains: it covers key areas of CS with finer subfields for specificity. (2) Diverse task forms: questions vary in format to simulate diverse real-world user queries. (3) CS-specific reasoning: it evaluates CS logical and arithmetic reasoning in addition to CS knowledge. (4) Multilinguality support: it supports assesses LLMs’ performance in different language environments. Based on these criteria, CS-Bench focuses on quadrilingual evaluation in English, Chinese, French, and German, covering four domains: Data Structure and Algorithm (DSA), Computer Organization (CO), Computer Network (CN), and Operating System (OS). Twenty-six fine-grained subfields, diverse task forms, and divisions of knowledge and reasoning are further developed to enrich the dimensions of assessment and simulate real-world scenarios.
@@ 2.2 Data Collection
Table 1: Comparison of perplexity (PPL) across evaluation datasets. The PPL of English and Chinese datasets is calculated on Llama2-7B and Qwen1.5-7B. “MC” denotes multiple-choice, and “ALL” denotes all tasks.
@@
TruthfulQA (MC) 7.73 C-Eval 11.47
MMLU 9.54 CMMLU 13.62
@@
Diverse data sources are key to achieving the sample diversity of CS-Bench. Our raw data originates from three sources: (1) Computer science-related questions obtained from publicly available online channels, such as professional exams and practice tests. (2)Knowledge-type questions obtained through the initial manual extraction and subsequent adaptation of blog articles from various computer-related websites. (3) Construction of teaching materials and examination papers authorized by the authors’ institutions.  [^1] More collection and processing procedures can be found in Appendix C. The latter two categories constitute the vast majority (over 70%) of the data, and these data are not directly exposed on the internet, effectively reducing the likelihood of LLMs encountering these questions during pre-training. We compare the perplexity (Jelinek et al., 1977) of models on CS-Bench English and Chinese datasets with several other prominent evaluation datasets in Table 1. In both English and Chinese, the perplexity of CS-Bench is comparable to or even higher than that of other datasets, further indicating the high quality of CS-Bench samples and the rarity of data leakage instances.
@@
The data processing relies on a team composed of five members, each holding a bachelor’s degree in computer science and receiving appropriate compensation. Initially, we parse questions and answers for each sample from the data sources either automatically or manually. Subsequently, we manually label questions with knowledge-type or reasoning-type tags depending on whether in-depth reasoning and calculation are required. For reasoning-type questions, we attempt to collect explanations from the data sources whenever possible; otherwise, we handle them through cross-annotation and verification among team members. We first construct Chinese data, then translate it into other languages using GPT-4, supplemented by manual checks, to create t multilingual data. Finally, we conduct thorough manual checks on the entire dataset to ensure quality. As this benchmark pertains to objective knowledge and reasoning in the field of computer science, the annotation content is not influenced by regional or cultural differences among annotators.
@@
CS-Bench is an evaluation benchmark supporting quadrilingual assessment, encompassing a total of 26 subfields across 4 domains, with a cumulative total of 9676 samples. These samples encompass various task formats including multiple-choice, assertion, fill-in-the-blank, and open-ended questions. Besides, CS-Bench assesses both knowledge-type and higher-order reasoning-type questions, with each reasoning question accompanied by an explanation. To validate the effectiveness of models, we randomly sample 10% of the data for validation, using the remaining 90% for testing. The statistics of CS-Bench are shown in Figure 1, with detailed exposition provided in Appendix C.
@@ 3.1 Experimental Setup
Due to the diverse task formats in CS-Bench, we first design question templates for each task type. For comprehension tasks (MC and Assertion), we use regex to match LLM’s predictions and then calculate their accuracy against the ground-truth answers. For generation tasks (FITB and Open-ended), due to the diversity of ground-truth answers, we score LLM’s predictions by GPT-4 using standard answers in CS-Bench as references. In detail, FITB questions are scored as either 0 or 1, while the score range for Open-ended questions is 1-10, which is then linearly mapped to a range of 0.1 to 1. Finally, scores are weighted based on the quantity of each type to derive the ultimate overall score. It is worth emphasizing that while using GPT-4 for scoring generation tasks may introduce a certain threshold for evaluation, its primary purpose is to simulate diverse task formats in real-world scenarios. Therefore, we encourage isolating comprehension tasks from CS-Bench to facilitate automatic evaluation with no need for GPT-4. We provide the details of the evaluation setup in Appendix D, where we also verify the validity of GPT-4 scoring through its consistency with manually scored results, and the effect of different scoring models.
@@
For open-source models, we selected Gemma-2B/7B (Team et al., 2024), Llama2-7B/13B/70B (Touvron et al., 2023), Llama3-8B/70B (Meta, 2024), Llama3.1-405B (Dubey et al., 2024), ChatGLM3-6B (THUDM, 2023), Baichuan2 (v2.0)-7B/13B (Yang et al., 2023), InternLM2-7B/20B (Cai et al., 2024) , Qwen1.5-4B/7B/14B/72B/110B (Alibaba, 2024), Mistral-7B (v0.2) (Jiang et al., 2023), Mixtral-8$\times$7B (v0.1) (Jiang et al., 2024), Mistral-Large-123B (MistralAI, 2024), and DeepSeekLLM-7B/67B (Bi et al., 2024). For closed-source commercial models, we utilized PaLM-2 (palm-2-chat-bison) (Anil et al., 2023), Claude-2.1 (Anthropic, 2023), Claude-3 (opus) (Anthropic, 2024a), Claude-3.5 (Sonnet) (Anthropic, 2024b), as well as GPT-3.5, GPT-4 (0125 version) (Achiam et al., 2023), GPT-4o (OpenAI, 2024a) and OpenAI-o1-mini/preview (0912 version) (OpenAI, 2024b) [^2] Due to the unique reasoning form of OpenAI-o1, we exclude it from Section 3.2 analyze it in Section 3.3.. To ensure the instruction-following abilities, we employ the official chat or instruction-tuned versions for all models. Details on these models are provided in Appendix D.4.
@@ 3.2 Main Results
Table 2 presents the overall results of all foundation models directly answering questions under the zero-shot setting  [^3] Due to space constraints, the results and analyses in other languages are provided in Appendix E.3.. In summary, the overall scores of models range from 39.86% to 72.29%, demonstrating CS-Bench’s effectiveness in distinguishing between the abilities of various models in the field of CS while also posing significant challenges to the best-performing existing models. Subsequently, we conduct a comprehensive analysis of the experimental results from various aspects including Foundation Models, Knowledge & Reasoning, Domains, and Task Formats.
Type Model Score Reasoning / Completion Tokens
Knowledge GPT-4o 76.95 - / 8.67
OpenAI-o1-mini 77.60 (+0.65) 251.0 / 269.06 ($\times$31.03)
OpenAI-o1-preview 83.61 (+6.66) 518.49 / 545.37 ($\times$62.9)
Reasoning GPT-4o 64.15 - / 56.43
OpenAI-o1-mini 76.12 (+11.97) 500.1 / 546.78 ($\times$9.69)
OpenAI-o1-preview 80.98 (+16.83) 1106.43 / 1198.78 ($\times$21.24)
Overall GPT-4o 72.29 - / 26.08
OpenAI-o1-mini 77.06 (+4.77) 341.8 / 370.29 ($\times$14.2)
OpenAI-o1-preview 82.65 (+10.36) 732.79 / 783.54 ($\times$30.04)
Figure 4: Comparison between OpenAI-o1 and GPT-4o. Reasoning Tokens refer to tokens used for internal reasoning in OpenAI-o1; Completion Tokens refer to all tokens consumed both internally and in external output. “$+$” and “$\times$” indicate performance gains and token consumption of OpenAI-o1 over GPT-4o.
Figure 5: The relationship between CS-Bench scores and the number of tokens consumed.
@@
Firstly, the closed-source models Claude3.5/GPT-4o represent the highest standard on CS-Bench, with a proficiency exceeding 70%. Secondly, leading open-source models (Mistral-Large-123B and Llama3.1-405B) have significantly narrowed the gap between open-source and closed-source models. On one hand, both models exhibit a substantial increase in parameters; on the other hand, current closed-source models are tending towards a balance between performance and efficiency, exemplified by GPT-4o-mini and Gemini Flash. Thirdly, newer models demonstrate significant improvements compared to earlier versions. For example, among models with scales below 10B, Llama3-8B performs the best, rivaling previous much larger-scale models and even surpassing Llama2-70B, indicating significant potential for compression in model parameters (Delétang et al., 2024). Lastly, while performance variations exist among models of different families at the same scale, models within the same family continue to improve with increasing scale on CS-Bench (see detailed scale analysis in Section 3.4).
@@
Overall, all models perform worse on reasoning (average 45.60%) compared to knowledge scores (average 60.61%), indicating that reasoning poses a greater challenge to LLMs compared to knowledge. As shown in Figure 3, there is a strong positive correlation between reasoning scores and knowledge scores. However, this correlation is not absolute. For instance, PaLM-2 has a higher knowledge score but a lower reasoning score compared to Claude-2.1, showing PaLM-2’s weakness in applying knowledge. Furthermore, more powerful LLMs demonstrate a stronger ability to use knowledge for reasoning compared to weaker LLMs. This is reflected in the much lower reasoning scores of weaker models relative to their knowledge scores. However, as the model’s capability increases, the growth in reasoning scores is more pronounced than that of knowledge scores, gradually bridging the gap between knowledge and reasoning abilities.
@@
As shown in Figure 3 (b) and Table 14, given the varying initial random scores, LLMs generally performs best on Assertion questions (average 63.14% across all models), followed by MC questions (average 55.45%), Open-ended questions (average 49.3%), and performs worst on FITB questions (average 41.58%). However, the variation in task format sensitivity is highly pronounced in weaker models, while stronger models can mitigate the disparities caused by different task formats, exhibiting robustness. For instance, Llama2-7B scores only 26.19% on Open-ended reasoning but 60.61% on Assertion reasoning, whereas GPT-4 scores comparably on both Open-ended reasoning (68.94%) and Assertion reasoning (67.68%).
3.3 OpenAI-O1 Result
Compared to general models’ reasoning processes, OpenAI-o1 models think before they answer, forming a long internal chain of thought through reasoning tokens. As shown in Figure 5, OpenAI-o1 significantly improves reasoning performance. While OpenAI-o1-mini nearly matches GPT-4o on knowledge-type questions, it boosts scores by 11.97 points on reasoning problems. Furthermore, OpenAI-o1-preview achieves substantial progress in both knowledge and reasoning questions. However, this performance gain comes at the expense of a high token consumption. For instance, OpenAI-o1-preview consumes an average of nearly 1.2K tokens on reasoning questions, 21.24 times that of GPT-4o. Figure 5 visualizes the relationship between model performance and completion token consumption. For reasoning questions, performance roughly increases logarithmically with the number of tokens, while the gains from token expenditure are lower for knowledge-type questions. We conduct a case study in Appendix E.7, comparing the responses of GPT-4o and OpenAI-o1.